# DAPPER: Discriminability-Aware Policy-to-Policy Preference-Based Reinforcement Learning for Query-Efficient Robot Skill Acquisition

### v1

@@ Abstract
Index Terms:
Reinforcement Learning, Human Factors and Human-in-the-Loop, Legged Robots.
@@ I Introduction
To address this issue, we focus on the ease of labeling and introduce preference discriminability as a key metric. This metric quantifies whether a human can determine how much two trajectories in a query are closer to their ideal behavior. When one trajectory is clearly closer to the ideal, labeling is straightforward. However, if both trajectories are similarly distant from the ideal, it becomes difficult for the human to judge, and a preference label may not be provided. By generating queries with high preference discriminability, we present clearer differences between trajectories, making them easier to label and thereby improving query efficiency, as a greater number of discriminable queries can be used to learn the human preference.
@@ II-B Improving Query Efficiency in PbRL
Ours: While previous approaches improve query efficiency by leveraging model confidence [2], addressing annotation noise [5], or actively selecting informative queries [6], they do not explicitly account for the labeling difficulty experienced by human annotators. In contrast, our framework improves query efficiency by modeling and leveraging human preference discriminability, which helps reduce indistinguishable queries and enhances query efficiency. This is particularly important in tasks where distinguishing task features is inherently difficult. If a query is not distinguishable to the human, a preference label cannot be obtained.
@@ III-A Reinforcement Learning
The most common task in reinforcement learning is to learn a policy $\pi:S\rightarrow A$ that prescribes the optimal action for the agent in each state. More specifically, the objective is often defined as maximizing the total rewards: $V_{\pi}(s)=\mathbb{E}_{\pi,\mathcal{P}}\Bigl{[}\textstyle\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{t})\,\Bigl{|}\,s_{0}=s\Bigr{]}$. Accordingly, common approaches have focused on learning an optimal policy by maximizing the value function $V$. Among various policy optimization algorithms, Proximal Policy Optimization (PPO) is recognized as an effective means to achieve this objective [12].
@@ IV Proposed Framework
while $i=1,2,...,\text{until converge}$ do
@@
            Update policy $\pi_{i}$ by RL update scheme (e.g., [13]) using $\mathcal{B}_{i}$
@@
$N$ Query number per iteration $10$
@@ IV-E3 Discriminability Reward
Using the dataset with inseparable labels $y\in\{0.5,\text{other}\}$ from the query dataset $\mathcal{D}$, the discriminator $D$ is updated. The corresponding loss function is defined as follows:
$\mathcal{L}_{\psi}$ $=$ $-\mathbb{E}_{(\pi_{i},\pi_{x},y)\sim\mathcal{D}}\Big{[}(1-\delta(y))\log D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))$ (8)
$+\delta(y)\log D(\mathcal{G}(\pi_{i}),\mathcal{G}(\pi_{x}))\Big{]}$
where $\delta$ is a delta function indicating discriminability, defined as $\delta=1\text{ if }y=0.5,\;\delta=0\text{ if }y=0\ \text{or}\ 1$.
@@
V Experiments
  • How does hyperparameter selection affect query efficiency and preference discriminability in our framework? (Section V-E)
  • Can the proposed method maintain its effectiveness as the number of feature dimensions that parametrize the target robot behaviors increases? (Section V-F)
  • Is our framework applicable to sim-to-real policy transfer? (Section V-G)
@@ V-A Common Settings
[…]
@@ V-A3 Labeling Queries from Simulated Annotator Model
[…]
@@ V-C1 Settings
This paper proposes a framework for generating queries that are easier for humans to distinguish, based on preference discriminability. This section examines whether actual humans can distinguish similar queries and how their ability declines with increasing similarity. To control for individual preference variability, we define a target feature $f^{*}$ and treat the policy closer to $f^{*}$ as the correct answer. We then analyze the proportion of indistinguishable responses ($y=0.5$) as a function of $d^{\text{disc}}$ to quantify actual-human preferecne discriminability. Additionally, we evaluate the accuracy of query labels when participants express a clear preference ($y\neq 0.5$). This experiment uses the “Posture” feature values defined in Table III.
@@ V-C2 Results
[…]
Figure 4: Evaluation of query efficiency: (a) Comparison to previous works. Top: Discrimination rate of queries. Middle: Converge query number for reaching the feature error threshold ($d^{\text{pref}}<0.02$), empirically defined by the convergence values of each method. If not converged within $2\times 10^{3}$ queries, the value is clipped at $2\times 10^{3}$. The threshold is shown as a red dashed line in the bottom plot. Bottom: Minimum values of feature error $d^{\text{pref}}$ achieved by the learned policies. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve shows the mean and variance over five experiments. (b) Learned policy features per query number. This experiments use $d^{\text{disc}}$ = “Medium”. The target feature $f^{*}$ is located at the intersection of the red dashed lines. The number of plotted points corresponds to the queries collected until the feature error falls below the threshold ($d^{\text{pref}}<0.02$), indicated by the red frame.
Figure 5: Performance comparison of DAPPER components: (Left) different temperature patterns of query sampling. (Right) different scaling coefficient patterns of policy exploration. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve plots the mean and variance over five experiments.
@@ V-D2 Results
As shown in Fig. 4(b), Baseline suffers from policy bias, where exploration is restricted by features of previously learned policies, resulting in slow progress toward the target features. In contrast, Ours retrains policies from scratch at each reward model update, avoiding policy bias and rapidly acquiring new features. Consequently, Ours outperforms the Baseline in terms of query efficiency.
V-E Influence of Different Query Exploration/Sampling
@@ V-E2 Results
Figure 6: Comparison of query efficiency in various tasks: “Trot only” means using feature extraction of trot patterns, “Normal” adds feature extraction of body posture as shown in Table III. This experiments use preference discriminability $d^{\text{disc}}$ = “Medium”, meaning that if a policy’s features are closer than a specified threshold, the policy is considered non-identifiable ($y=0.5$). Each curve plots the mean and variance over five experiments.

### v2

@@ I Introduction
Legged locomotion has received considerable attention, with growing interest in enabling robots to acquire walking behaviors that reflect human preferences rather than relying solely on task-specific objectives [1, 2]. Existing approaches mainly rely on reinforcement learning [3, 4] or imitation learning [5, 2]. Reinforcement learning requires task-specific reward parameter tuning that depends heavily on expert knowledge, while imitation learning relies on demonstrations that demand operator proficiency with robot interfaces and motion strategies. Both introduce substantial costs for users, making it difficult to realize locomotion behaviors aligned with individual preferences. In this paper, we focus on quadruped locomotion and propose a preference-based learning framework that reduces user effort while shaping policies according to human preferences.
@@
To address this issue, we focus on the ease of labeling and introduce preference discriminability as a key metric. This metric quantifies whether a human can determine how much two trajectories in a query are closer to their ideal behavior. When one trajectory is clearly closer to the ideal, labeling is straightforward. However, when both trajectories are similarly undesirable or equally good, humans may struggle to provide a clear preference, not because the trajectories themselves are indistinguishable, but because their underlying preferences are indistinguishable. By generating queries with high preference discriminability, we present clearer differences between trajectories, making them easier to label and thereby improving query efficiency, as a greater number of discriminable queries can be used to learn the human preference.
@@ II-B Improving Query Efficiency in PbRL
Active Reward Learning: This approach collects queries from a pre-generated trajectory dataset and trains a policy only once using the reward function learned from those queries. Following information-theoretic objective, it implicitly favors trajectory pairs that are easier for humans to distinguish, since indistinguishable queries yield little information gain. However, because it depends entirely on a fixed dataset of candidate trajectories, it assumes that optimal and easily comparable pairs already exist in the dataset. Without a mechanism to generate new trajectories through policy optimization, it cannot expand behavioral variations for more discriminable queries [14, 10].
Reward Re-calculation: This approach improves query efficiency by recalculating the reward values of previously collected trajectories, enabling the reuse of past dataset without requiring additional queries. The agent can effectively expand its training dataset and refine the reward model more rapidly. This helps reduce the reliance on new annotations and contributes to more stable policy learning [15, 10].
Ours: While previous approaches improve query efficiency by leveraging model confidence [7], addressing annotation noise [10], actively selecting informative queries [11], selecting queries based on information gain [14], or reusing past data through reward re-calculation [15], they do not explicitly account for the labeling difficulty experienced by human annotators. In contrast, our framework improves query efficiency by modeling and leveraging human preference discriminability, which helps reduce indistinguishable queries and enhances query efficiency. This is particularly important in tasks where distinguishing task features is inherently difficult. If a query is not distinguishable to the human, a preference label cannot be obtained.
@@ III-A Reinforcement Learning
The most common task in reinforcement learning is to learn a policy $\pi:S\rightarrow A$ that prescribes the optimal action for the agent in each state. More specifically, the objective is often defined as maximizing the total rewards: $V_{\pi}(s)=\mathbb{E}_{\pi,\mathcal{P}}\Bigl[\textstyle\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{t})\,\Bigl|\,s_{0}=s\Bigr]$. Accordingly, common approaches have focused on learning an optimal policy by maximizing the value function $V$. Among various policy optimization algorithms, Proximal Policy Optimization (PPO) is recognized as an effective means to achieve this objective [3, 4].
@@ IV Proposed Framework
while i=1,2,,until convergei=1,2,\ $\text${until converge} do
@@
    Update $\pi_{i}$ by RL update scheme (e.g. [3]) using $\mathcal{B}_{i}$
@@
NN Query number per iteration $10$
@@ IV-E3 Discriminability Reward
The discriminator $D$ is trained on the query dataset $\mathcal{D}$, where the label $y$ can take inseparable values $y\in\{0.5,\text{other}\}$. The loss function is defined as follows:
$\mathcal{L}_{\psi}$ $=$ $-\mathbb{E}_{(\pi_{i},\pi_{x},y)\sim\mathcal{D}}\Big[y\log D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))$ (8)
$\quad+\;(1-y)\log\big(1-D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))\big)\Big],$
where $D$ is defined as a symmetric function, i.e., $D(\pi_{x},\pi_{i})=D(\pi_{i},\pi_{x})$. This formulation is equivalent to the standard binary cross-entropy loss, encouraging the discriminator $D$ to assign high probability when two policy embeddings are distinguishable, and low probability otherwise.
@@
V Experiments
  • How does our framework perform when the threshold of preference discriminability is set substantially above or below the range observed in actual human responses? (Section V-E)
  • How does our framework affect computational time efficiency? (Section V-F)
  • How does hyperparameter selection affect query efficiency and preference discriminability in our framework? (Section V-G)
  • Can our framework effectively learn from VLM-based preference labels, which rely on world knowledge aggregated from large-scale data, when used as substitutes for human queries that may include labeling errors or indistinguishable cases? (Section V-H)
  • Can the proposed method maintain its effectiveness as the number of feature dimensions that parametrize the target robot behaviors increases? (Section V-I)
  • Is our framework applicable to sim-to-real policy transfer? (Section V-J)
Figure 3: Analysis of actual human query responses for determining the threshold of preference discriminability $d^{\text{disc}}$ in the simulated annotator model: A questionnaire was conducted using a console system, as shown in (b). At the top of the screen, a reference video was displayed, showing a policy that matched a predefined target behavior. Below the reference video, two candidate policy videos were shown for comparison. Participants were asked to choose the trajectory closer to the reference by selecting one of three options: “Left is better”, “Right is better”, or “Can’t decide”. The candidate policies were prepared by exhaustively combining the listed values of $\text{Body Height [cm]}\in\{29,37,45,53,61\}$ and $\text{Body Incline Angle [degree]}\in\{-8,-4,0,4,8\}$. For each combination of these parameters, a policy was trained using DAPPER under an idealized setting where $d^{\text{disc}}=0$, by optimizing the behavior to match the specified Body Height and Body Incline Angle values until convergence. Participants were instructed to compare only two features defined by $f^{*}$: Body Height and Body Incline Angle. Gait patterns were excluded. They could prioritize the features freely, but were told to consider both. Policy pairs were sampled from (a), and 105 predefined queries were presented sequentially. The queries were selected to ensure diversity in preference discriminability $d^{\text{disc}}$. All participants evaluated the same set of queries, and the order of the queries was randomized for each participant. The results shown in (c) represent the mean and standard deviation across five participants.
@@ V-A Common Settings
All experiments are conducted using the ANYmal quadrupedal robot and the legged_gym RL-framework based on the IsaacGym physics simulator [1]. The trained policy is evaluated at $50\text{\,}\mathrm{H}\mathrm{z}$ predicting target joint positions, which are tracked by per joint PID controller running at $400\text{\,}\mathrm{H}\mathrm{z}$. In all experiments, the learning parameter settings follow those of previous works, including simulation parameters, specifically state-action definition and network architectures [1]. Policy learning utilized Lagrangian Proximal Policy Optimization (LPPO) [3, 4] with parallel environments, and all experiments are conducted using a Nvidia GeForce RTX 4090 GPU. To stabilize the $R^{D}$ signal particularly during the early training phase, we introduced Monte Carlo dropout [16] into the discriminator $D$. Furthermore, to prevent the discriminator $D$ from overfitting to earlier policies when continuously updated across iterations, their parameters were reset and retrained from scratch at each iteration.
@@ V-A3 Labeling Queries from Simulated Annotator Model
Figure 4: Evaluation of query efficiency: (a) Comparison to previous works. Top: Discrimination rate of queries. Middle: Converge query number for reaching the feature error threshold ($d^{\text{pref}}<0.02$), empirically defined by the convergence values of each method. If not converged within $2\times 10^{3}$ queries, the value is clipped at $2\times 10^{3}$. The threshold is shown as a red dashed line in the bottom plot. Bottom: Minimum values of feature error $d^{\text{pref}}$ achieved by the learned policies. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve shows the mean and standard deviation over five experiments. (b) Learned policy features per query number. This experiments use $d^{\text{disc}}$ = “Medium”. The target feature $f^{*}$ is located at the intersection of the red dashed lines. The number of plotted points corresponds to the queries collected until the feature error falls below the threshold ($d^{\text{pref}}<0.02$), indicated by the red frame.
@@ V-B Comparing Methods
  • PEBBLE [15]: Extends the Baseline with unsupervised pre-training and off-policy learning. The agent explores with intrinsic rewards maximizing state entropy to generate diverse behaviors before learing from queries. Past reward values are re-calculated and utilized after each reward model update, enabling sample-efficient policy learning.
@@ V-C1 Settings
This paper proposes a framework for generating queries that are easier for humans to distinguish, based on preference discriminability. This section examines whether actual humans can distinguish similar queries and how their ability declines with increasing similarity. To control for individual preference variability, we define a target feature $f^{*}$ and treat the policy closer to $f^{*}$ as the correct answer. We then analyze the proportion of indistinguishable responses ($y=0.5$) as a function of $d^{\text{disc}}$ to quantify actual-human preference discriminability. Additionally, we evaluate the accuracy of query labels when participants express a clear preference ($y\neq 0.5$). This experiment uses the “Posture” feature values defined in Table III.
@@ V-C2 Results
The experimental results are shown in Fig. 3. As seen in Fig. 3(c), when $d^{\text{disc}}$ is low, around $60\text{\,}\%$ of queries were judged as indistinguishable. As $d^{\text{disc}}$ increases, distinguishability improves, reaching $100\text{\,}\%$ for all participants when $d^{\text{disc}}>0.4$. Label accuracy for distinguishable queries exceeded $80\text{\,}\%$ at all $d^{\text{disc}}$ levels, and reached $100\text{\,}\%$ when $d^{\text{disc}}>0.3$. These results suggest that $d^{\text{disc}}>0.4$ ensures consistent discrimination, while lower values depend on individual ability. Based on this, we define three threshold levels as criteria for automatic labeling in the subsequent sections: “Small” ($0.2$), “Medium” ($0.3$), and “Large” ($0.4$).
Figure 5: Performance comparison across different preference discriminability thresholds: The convergence query number denotes the point at which the learned policy reaches the feature error threshold ($d^{\text{pref}}<0.02$). If convergence is not achieved within $2\times 10^{3}$ queries, the value is clipped at $2\times 10^{3}$. Each curve shows the mean and standard deviation over five experiments.
@@ V-D2 Results
As shown in Fig. 4(b), the Baseline suffers from policy bias that constrains exploration to the features of previously learned policies and slows progress toward the target features $f^{*}$. In contrast, Ours w/o $R^{D}$ initially learns policies with similar feature representations because it lacks $R^{D}$, but retraining policies from scratch at each reward model update reduces the influence of policy bias and occasionally allows larger jumps to new feature regions. However, without the discriminability reward $R^{D}$ to actively encourage exploration, the policy tends to converge to local optima due to lack of discriminable queries. Meanwhile, Ours (i.e., with $R^{D}$) integrates training-from-scratch with the discriminability reward, consistently acquiring diverse and distinguishable policies that successfully converge toward the target feature value $f^{*}$.
Figure 6: Analysis of computational time: (Left) Total computational time per iteration $i$. As shown in Fig. 2 and Algorithm 1, each iteration $i$ consists of three processes: “1) Policy Learning”, “2) Query Collection”, and “3) Reward Update”. In this plot, the stacked bar represents the total computational time, computed as the sum of the mean time spent on each process (averaged across the five trials). Note that the “3) Reward Update” process takes only a few seconds and is barely visible in the plot. (Right) Total computational time until convergence, defined as the point where the learned policy’s feature error satisfies the threshold ($d^{\text{pref}}<0.02$). The bars show the mean time to converge, and the error bars denote the standard deviation across the five trials. As shown in Algorithm 1, the parameter $J$ represents the number of policy learning iterations used to train $\pi_{i}$ in “1) Policy Learning” within each cycle. The computation time for “2) Query Collection” is estimated at about 70 seconds based on Fig. 3, where one human response took around 7 seconds on average and each iteration included $N=10$ queries.
V-E Analysis of Sensitivity to Preference Discriminability Threshold
@@ V-E1 Settings
As shown in Fig. 4, our framework demonstrates improved query efficiency under preference discriminability thresholds corresponding to the range observed in actual human responses. In this section, we investigate how the framework behaves when the threshold is set beyond this human-observed range, either higher or lower, to observe its behavior under more challenging or relaxed discriminability conditions. For comparison, the Baseline is used to examine how our framework behaves differently under these varying threshold conditions. This experiment uses the target feature value of “Posture”.
V-E2 Results
As shown in Fig. 5, Ours consistently outperforms the Baseline across a wide and fine-grained range of thresholds. The results also show that the performance gap narrows when the preference discriminability threshold is very low. However, even when the threshold is set to zero, meaning that all queries are distinguishable, the Baseline still performs worse than Ours because policy bias prevents it from learning substantially different behaviors within a small number of iterations $i$. Conversely, as the threshold increases and discriminable queries become scarce, policy learning becomes increasingly difficult, and Ours ultimately fails to learn policies.
V-F Analysis of Computational Time Efficiency
V-F1 Settings
As shown in Fig. 4, this paper demonstrates improvements in preference discriminability and the resulting query efficiency. However, achieving these improvements introduces additional computational components. Specifically, a new policy is retrained from scratch at each iteration $i$ to avoid policy bias, and a discriminator $D$ is introduced as an additional component to estimate preference discriminability. In this experiment, we evaluate whether the gain in query efficiency achieved by our framework outweighs the additional computational cost in terms of total computation time until policy-learning converge, by examining whether it remains more time-efficient than the Baseline, which does not include these additional components. We evaluate the Baseline under different numbers of policy learning iterations ($J\in\{300,30,10\}$), since the Baseline does not initialize the policy at each iteration $i$ and can continue learning with fewer policy learning steps. This experiment uses target feature values of “Posture” under the “Medium” preference discriminability condition.
V-F2 Results
The results are shown in Fig. 6. Regarding total time per iteration $i$, the “1) Policy Learning” process takes about 12–13 minutes, accounting for most of the total 13–14 minutes. In contrast, the “3) Reward Update” requires only a few seconds, having a negligible effect on the total time. Reducing the number of policy learning iterations $J$ naturally decreases the overall computation time per iteration $i$. Regarding the total time until convergence, with $J=30$, the Baseline reduces the overall time to roughly $40\text{\,}\%$ of that with $J=300$, but it is still approximately twice as long as Ours. However, when $J$ is reduced to $10$, the time for “1) Policy Learning” becomes smaller, but “2) Query Collection” dominates, and more iterations are required for convergence, resulting in longer total training time. Overall, even when accounting for retraining from scratch and the additional discriminator $D$, Ours demonstrates superior time efficiency compared to the Baseline.
V-G Influence of Different Query Exploration/Sampling
V-G1 Settings
@@ V-G2 Results
Figure 7: Performance comparison of DAPPER components: (Left) different temperature patterns of query sampling. (Right) different scaling coefficient patterns of policy exploration. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve plots the mean and standard deviation over five experiments.
Figure 8: Prompt for VLM-based preference labeling: We used OpenAI GPT-5 model with the instruction text shown in the figure to generate preference labels. Three videos were used in each comparison: two candidate videos and one reference video. Each video is 3 seconds long at 30 Hz, from which three frames were sampled at 0.5, 1.5, and 2.5 seconds and downscaled from the original resolution of 1,600×900 pixels (the same resolution as in Fig. 3) to 800×450 pixels to reduce computational and resource costs. These processed frames, together with the textual instructions, were used as inputs. The model was then asked to output one of three labels: “Left is better”, “Right is better”, or “Can’t decide”.
V-H Applicability of VLM-Based Preference Labels as Human Query Substitutes
V-H1 Settings
Vision-language models (VLMs) provide an attractive alternative to human annotators for generating preference labels [19, 20]. Using VLMs instead of humans could significantly reduce annotation cost, but it also introduces challenges such as noisy or overconfident predictions. Highlighting this alternative is particularly relevant for PbRL, where both sample efficiency and robustness to noisy labels are key for practical deployment. In this experiment, we conduct a preliminary study to examine whether our framework can leverage VLM-based preferences effectively. The prompt and configuration provided to the VLM are shown in Fig. 8. While VLMs offer a potential substitute for human preference annotations, they often tend to be overconfident when choosing between options and rarely select the “Can’t decide” label. This behavior may stem from their underlying training or preference-oriented fine-tuning. To address this, as illustrated in Fig. 8, we assumed that the VLM’s frame-wise estimates are independent and identically distributed and applied Welch’s t-test. If no statistically significant difference was found, the query was mapped to the “Can’t decide” label. This experiment used the target feature value of “Posture” as defined in Table III.
V-H2 Results
As shown in Table IV, VLM-based preference labels exhibit both discriminability and accuracy limitations. The Baseline achieved only about $60\text{\,}\%$ label accuracy, which is extremely low considering that the labels are binary. Moreover, nearly half of the queries could not be assigned a label, leading to almost no learning in terms of the feature error $d^{\text{pref}}$. These results indicate that VLMs tend to generate overconfident outputs, avoiding the “Can’t decide” option and thus producing unreliable supervision. In contrast, our framework leverages preference discriminability to reduce the number of indistinguishable queries, improving label accuracy by approximately $20\text{\,}\%$ over the Baseline. Consequently, the learned policy reached the feature error threshold ($d^{\text{pref}}<0.02$) defined in Fig. 4, demonstrating that useful signals can be extracted from VLM-generated labels when query design accounts for discriminability.
V-I Scalability with Respect to Feature Dimensions
V-I1 Settings
V-I2 Results
TABLE IV: Performance evaluation with VLM-based preference labeling: Each trial was conducted with up to 1,000 queries. Each value represents the mean and standard deviation over five experiments.
Evaluation Metric Ours Baseline
Indiscriminable Rate $\%$] $23.6\pm 15.7$ $46.9\pm 16.0$
Label Accuracy $\%$] $84.7\pm 12.8$ $61.9\pm 15.2$
Minimum Feature Error $d^{\text{pref}}$ $0.005\pm 0.002$ $0.318\pm 0.061$
V-J Real-World Application
In this section, we apply the learned control policies from simulation to real-world environments using a sim-to-real transfer approach [1]. The transfer policy is trained in a domain-randomized environment over sensor noise and physical parameters, using well-validated parameter sets from previous work [1]. During policy learning, the feature representation $f$ used as input to the learned preference reward model $R^{H}$ is treated as privileged information available only in simulation for reward computation, and therefore domain randomization is not applied to $f$. As shown in Fig. 10, the learned policies were applied in the real world. This validation follows the same sim-to-real setup as in previous work [1], without additional tunings. For details of the robot motion, please refer to the supplementary videos.
Figure 9: Comparison of query efficiency in various tasks: “Trot only” means using feature extraction of trot patterns, “Normal” adds feature extraction of body posture as shown in Table III. This experiments use preference discriminability $d^{\text{disc}}$ = “Medium”, meaning that if a policy’s features are closer than a specified threshold, the policy is considered non-identifiable ($y=0.5$). Each curve plots the mean and standard deviation over five experiments.